------------------------------------------------------------ Date: Fri Aug 3 08:30:58 2001, Shift: 00:00-08:00 Supervisor: KH, with: Mads Log book#7 pages 9 - 14 Shift Summary: ------------------------------------------------------------ 2 hrs 45 min of beam during this shift. When we had beam, it was with 56 bunches, so intensity was a factor of 10 greater than previously 3 data runs 3975 8 min 8000 collisions 3976 15 min 6123 collisions 3977 2 hr 20 min estimate ~100k collisions as per SuperMon BrNtuple not finished processing this run No detector problems Machine status: 0:00 - 4:00 Quench Detection FEC problem. Also Cryo problem 4:00 - 4:40 Inject, ramp and run 4:40 - 5:30 Yellow abort kicker problem 5:30 - 8:00 Inject, ramp and run 8:00 - 200 MHz RF development with 6 bunch operations ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Fri Aug 03 2001 - 08:31:50 EDT